# Heegaard Floer Surgery Formula and Cosmetic Surgeries

### v1

@@
Abstract.
Two Dehn surgeries on a knot are called cosmetic if they yield homeomorphic three-manifolds. We show for a certain family of null-homologous knots in any closed orientable three-manifold, if the knot admits cosmetic surgeries with a pair of positive surgery coefficients, then the coefficients are both greater than $1$. In addition, for this family of knots, we show that $1/q$ Dehn surgery for $q$ at least $2$ is not homeomorphic to the original three-manifold. The proofs of these results use the mapping cone formula for the Heegaard Floer homology of Dehn surgery in terms of the knot Floer homology of the knot; we provide a new proof of this formula for integer surgeries in $\operatorname{Spin}^{c}$ structures with nontorsion first Chern class.
1. Introduction
@@
Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that $Y-K$ is irreducible and not homeomorphic to the solid torus. If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of $Y-K$ which takes one slope to the other.
The existence of cosmetic surgeries has been ruled out in many cases for knots in $S^{3}$. For instance, Ozsváth and Szabó proved that if $K\subset S^{3}$ admits cosmetic rational surgeries, then either the resulting manifold is an $L$-space or the surgery slopes have opposite signs [OS10]. Much more is known when considering purely cosmetic surgeries. Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are ±$\pm 2$ or ±$\pm 1/q$ for some integer $q$ [Han19].
@@
Let $Y$ be a closed, oriented three-manifold, and $K\subset Y$ a null-homologous knot. We say that $K$ is blow-up unknotted if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $[G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
@@
The property of $K$ being blow-up unknotted implies that $K$ is null homotopic.
For example, by Kirby calculus, $K$ is a blow-up unknotted knot if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
@@
Let $K\subset Y$ be a nontrivial blow-up unknotted knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $g(K)\geq 2$, then $Y_{1}(K)\ncong\pm Y$.
@@
Acknowledgements The author would like to thank Yi Ni for advising him in this project.
2. Preliminaries on Heegaard Floer Homology
In this section, we establish some notation and collect some results that will be needed in later sections.
@@ 2. Preliminaries on Heegaard Floer Homology
Knot Floer homology associates to a knot $K$ in a closed, oriented three-manifold $Y$ a ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered ${\mathbb{Z}}[U]$-complex $C=CFK^{\infty}(Y,K)$. Write the two factors of the ${\mathbb{Z}}\oplus{\mathbb{Z}}$ filtration as $(i,j)$. Let $S$ be a region in the plane such that for all $(i,j)\in S$ and all $(i^{\prime},j^{\prime})\geq(i,j)$, we have that $(i^{\prime},j^{\prime})\in S$. Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$ as in Ozsváth-Szabó. There are two canonical chain maps $v_{s}^{+}:A_{s}^{+}\to B^{+}$ and $h_{s}^{+}:A_{s}^{+}\to B^{+}$, where $v_{s}^{+}$ is the vertical projection, and $h_{s}^{+}$ first projects $A_{s}^{+}$ to $C\{j\geq s\}$, then maps to $C\{j\geq 0\}$ via $U^{s}$, and finally maps to $B^{+}$ via a chain homotopy equivalence $C\{j\geq 0\}\to C\{i\geq 0\}$ [OS08].
Let $K\subset Y$ be null-homologous. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and let $n\in{\mathbb{Z}}$. The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the cobordism $W_{n}^{\prime}(K)$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\mod 2n.$
For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of
$CF^{+}(Y_{n}(K))$. Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
@@
${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\prod_{s\in{\mathbb{Z}}|s\equiv i\mod n}B_{\mathfrak{t}}^{+}$
@@
Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$ [OS08].
@@
Let $K\subset Y$ be a null-homologous knot and let $F$ be a Seifert surface for $K$ in the homology class $\phi$. Given $n>0$ and $t\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ be the natural two-handle cobordism, and let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping of $\partial F$ with a disk. Let $\widehat{\phi}_{n}=[\widehat{F}_{n}]\in H_{2}(W_{n}^{\prime}(K))$. Let $k$ be an integer satisfying $k\equiv t\mod n$ and $|k|\leq n/2$. Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{t}_{k}|Y=\mathfrak{n}_{k}|Y=\mathfrak{s}$ and
$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
and let $\mathfrak{s}_{t}=\mathfrak{t}_{k}|Y_{n}(K)=\mathfrak{n}_{k}|Y_{n}(K)$.
@@
When $n\geq 2g(F)$, the chain complex
$CF^{+}(Y_{n}(K),\mathfrak{s}_{t})$ is represented by the chain complex $A_{k}^{+}$, in the sense that there is an isomorphism
$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{s}_{t})\to A_{k}^{+}.$
Moreover, the following squares commute:
{tikzcd}
{tikzcd}
@@
${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\prod_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
@@
Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$ [OS10].
@@ 3. Integer Surgeries
Fix $n>0$. The long exact sequence in Theorem 2.2 decomposes over $\operatorname{Spin}^{c}$ structures on $Y$. Setting $m=nk$ for large $k$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e. given $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ where
(3.1) $f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\mod 2n(k+1),$
$\text{and }s\equiv i\mod n\},$
$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
A strengthening of the proof of Proposition 3.5 in [Ni20] shows the following:
@@
Let $p=\max\{k+g(F),-2k+g(F)\}$. Suppose that $n\geq\max\{p,2g(F)\}$, $\mathfrak{s}\in\mathcal{X}_{k}\setminus\{\mathfrak{x}_{k},\mathfrak{y}_{k}\},$ then the map $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through the map
@@
Proof.
The proof is exactly the same as in [Ni20], except we note that
$\frac{l(l+1)n}{2}-lk-g\geq n-p$
whenever $l\neq 0,-1$. From here, it follows that $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through $U^{n-p}$. ∎
@@
is identified with $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
@@
Recall that $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ as in Equation 3.1. Since $c_{1}(\mathfrak{t})$ is nontorsion, $H_{*}(A_{i}^{+})$ is a finitely generated abelian group. So $U^{p}|H_{*}(A_{i}^{+})=0$ when $p$ is greater than a constant $C_{1}$ independent of $n,k$. By Theorem 2.3, this implies that $U^{p}|HF^{+}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $p>C_{1}$, $n(k+1)>2g$.
Recall that we choose representatives
$-\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
@@
$l\leq\frac{n(k+1)}{2}+g\leq n(k+1)$
for large $k$. Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain equivalence
$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}.$

### v2

@@
Abstract
Two Dehn surgeries on a knot are called cosmetic if they yield homeomorphic three-manifolds. We show for a certain family of null-homologous knots in any closed orientable three-manifold, if the knot admits cosmetic surgeries with a pair of positive surgery coefficients, then the coefficients are both greater than $1$. In addition, for this family of knots, we show that +1/q+1/q Dehn surgery is not homeomorphic to the original three-manifold. The proofs of these results use the mapping cone formula for the Heegaard Floer homology of Dehn surgery in terms of the knot Floer homology of the knot; we provide a new proof of this formula for integer surgeries in $\operatorname{Spin}^{c}$ structures with nontorsion first Chern class.
Keywords: Heegaard Floer Homology, knot Floer homology, Dehn surgery, cosmetic surgery
1 Introduction
@@
Suppose $K$ is a knot in a closed oriented three-manifold $Y$ such that YKY-K is irreducible and not homeomorphic to the solid torus. If two different Dehn surgeries on $K$ are purely cosmetic, then there is a homeomorphism of YKY-K which takes one slope to the other.
The existence of cosmetic surgeries has been ruled out in many cases for knots in $S^{3}$. For instance, Ozsváth and Szabó proved that if $K\subset S^{3}$ admits cosmetic rational surgeries, then either the resulting manifold is an $L$-space or the surgery slopes have opposite signs [OS10]. Much more is known when considering purely cosmetic surgeries. Hanselman proved that any pair of purely cosmetic Dehn surgeries on knots in $S^{3}$ must have slopes that are $\pm 2$ or $\pm 1/q$ for some integer $q$ [HAN19].
@@
Let $Y$ be a closed, oriented three-manifold, and $K\subset Y$ a null-homologous knot. We say that $K$ is blow-up bounding if there exists a taut Seifert surface $F$ for $K$, some natural number $N$, and a properly-embedded smooth disk $G$ inside $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ with $\partial G=K\times\{0\}$ and $G]=\iota_{*}[F]\in H_{2}(X,K\times\{0\})$ for $\iota:(Y,K)\to(X,K\times\{0\})$ the inclusion.
We say that $K$ is blow-up unknotted if $K$ is obtained from the unknot in $Y$ by a sequence of crossing changes that change a positive crossing to a negative one.
By Kirby calculus, $K$ is blow-up bounding if $K$ is blow-up unknotted.
@@
The property of $K$ being blow-up bounding implies that $K$ is null homotopic.
@@
Let $K\subset Y$ be a nontrivial blow-up bounding knot and $q\geq 2$, then $Y_{1/q}(K)\ncong\pm Y$. Furthermore, if $K$ is a nontrivial blow-up unknotted knot, then $Y_{1}(K)\ncong\pm Y$.
@@
2 Preliminaries on Heegaard Floer Homology
In this section, we establish some notation and collect some results, most of which are copied from [OS08], that will be needed in later sections.
@@ 2 Preliminaries on Heegaard Floer Homology
[…]
Write the two factors of the $\mathbb{Z}$$\oplus{\mathbb{Z}}$ filtration as $(i,j)$. There is a $U$-action on $C$ that is an isomorphism of bifiltration (1,1)(-1,-1).
We will often use the following fact, for example see [HW18, Reduction Lemma].
Lemma 2.3.
With coefficients in any field $\mathbb{F}$}, the knot Floer complex $(C,\partial)$ for any knot $K\subset Y$ is $({\mathbb{Z}}\oplus{\mathbb{Z}})$-filtered chain homotopy equivalent to a chain complex $(\overline{C},\overline{\partial})$ such that the restriction of the differential $\overline{\partial}$ to any $\overline{C}\{i,j\}$ is 0 and $\overline{C}\{0,j\}$ is isomorphic to $\widehat{HFK}(K,j)$.
Knot Floer homology is known to detect the genus of the knot, and this is true for knots in general three-manifolds.
Theorem 2.4.
[NI10, Theorem 2.4] For any field $\mathbb{F}$}, if $K\subset Y$ is a knot of genus $g$, then
$\widehat{HFK}(Y,K,s;{\mathbb{F}})=0$
if $|s|>g$. Furthermore, $\widehat{HFK}(Y,K,g;{\mathbb{F}})\neq 0$.
Define $A_{s}^{+}=C\{\max(i,j-s)\geq 0\}$ and $B^{+}=C\{i\geq 0\}$; these are chain complexes because the differential decreases both filtrations. There are two canonical chain maps $v_{s}^{+}:A_{s}^{+}\to B^{+}$ and $h_{s}^{+}:A_{s}^{+}\to B^{+}$, where $v_{s}^{+}$ is the vertical projection, and $h_{s}^{+}$ first projects $A_{s}^{+}$ to $C\{j\geq s\}$, then maps to $C\{j\geq 0\}$ via $U^{s}$, and finally maps to $B^{+}$ via a chain homotopy equivalence $C\{j\geq 0\}\to C\{i\geq 0\}$ [OS08].
For all $s\in{\mathbb{Z}}$, there is a chain homotopy equivalence $\phi$ between $A_{s}^{+}$ and $A_{-s}^{+}$: first, the $U^{s}$ map is an isomorphism
$A_{s}^{+}=C\{\max(i,j-s)\geq 0\}\to C\{\max(i,j-s)\geq-s\},$
then the change of basepoint map that swaps the roles of $w$ and $z$ in the doubly-pointed Heegaard diagram is a chain homotopy equivalence
$C\{\max(i,j-s)\geq-s\}\to C\{\max(i,j+s)\geq 0\}=A_{-s}^{+}.$
Let $\phi$ be the composition of these two maps. It is clear from the definitions that the following lemma holds:
Lemma 2.5.
For all $s\in{\mathbb{Z}}$, $v_{s}^{+}$ and $h_{s}^{+}$ are chain homotopy equivalent to $h_{-s}^{+}\circ\phi$ and $v_{-s}^{+}\circ\phi$, respectively.
Since all maps involved are $U$-equivariant, the above lemma also holds for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$.
Lemma 2.6.
Let $g$ be the genus of $K$, then for all $\delta\geq 0$ and $s\geq g+\delta$, $v_{s}^{\delta}$, and $h_{-s}^{\delta}$ induce isomorphisms on homology, and $v_{-s-1}^{\delta}$ and $h_{s+1}^{\delta}$ are nullhomotopic.
Proof.
It is shown in the proof of the adjunction inequality in [OS04a, Theorem 5.1] that one can construct a special Heegaard diagram for $(Y,K)$ such that all generators of $\widehat{CFK}(Y,K)$ lie in Alexander grading greater than or equal to g-g. Thus, all generators for $A_{s}^{\delta}$ with $i\geq 0$ must have $j\geq-g$, from which it follows that $h_{-s}^{\delta}$ induces an isomorphism on homology and $v_{-s-1}^{\delta}$ is null-homotopic. By the $CF^{\delta}$ version of Lemma 2.5, we obtain the statements about $v_{s}^{\delta}$ and $h_{s+1}^{\delta}$. ∎
Let $K\subset Y$ be null-homologous. Let $F$ be a Seifert surface for $K$ and let $\hat{F}$ be the closed surface obtained by capping off $\partial F$ with a disk. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, which need not have torsion first Chern class, and let $n\in{\mathbb{Z}}$. The space of $\operatorname{Spin}^{c}$ structures over $Y_{n}(K)$ which are $\operatorname{Spin}^{c}$-cobordant to $\mathfrak{t}$ over the natural two-handle cobordism $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ is identified with ${\mathbb{Z}}/n{\mathbb{Z}}$ by $\mathfrak{u}\mapsto i$, where $\mathfrak{s}$ is an extension of $\mathfrak{u}$ over $W_{n}^{\prime}(K)$ with the property that
$\langle c_{1}(\mathfrak{s}),[\hat{F}]\rangle-n\equiv 2i\pmod{2n}$.
For $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ denote the corresponding summand of $CF^{+}(Y_{n}(K))$. Let $A_{s,\mathfrak{t}}^{+},B_{\mathfrak{t}}^{+}$ denote the corresponding chain complexes associated to the knot filtration on $CF^{+}(Y,\mathfrak{t})$ [OS08].
@@
${\mathbb{A}}_{i,\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}A_{s,\mathfrak{t}}^{+}\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{\mathfrak{t}}^{+}=\bigoplus_{s\in{\mathbb{Z}}|s\equiv i\pmod{n}}B_{\mathfrak{t}}^{+}$
@@
Let ${\mathbb{X}}_{i,\mathfrak{t}}^{+}(n)$ denote the mapping cone of $D_{n,i,\mathfrak{t}}^{+}$; we prove in the next section that this chain complex computes $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ [OS08].
@@
Let $K\subset Y$ be a null-homologous knot and let $F$ be a Seifert surface for $K$ in the homology class $\phi$. Given $n>0$ and $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, let $\widehat{F}_{n}\subset W_{n}^{\prime}(K)$ be the closed surface obtained by capping off $\partial F$ with a disk. Let $\widehat{\phi}_{n}=[\widehat{F}_{n}]\in H_{2}(W_{n}^{\prime}(K))$. Let $k$ be an integer satisfying $k\equiv i\pmod{n}$ and $|k|\leq n/2$. Let $\mathfrak{x}_{k},\mathfrak{y}_{k}\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ satisfy $\mathfrak{x}_{k}|Y=\mathfrak{y}_{k}|Y=\mathfrak{t}$ and
$\langle c_{1}(\mathfrak{x}_{k}),\hat{\phi}_{n}\rangle=2k-n,\hskip 14.22636pt\langle c_{1}(\mathfrak{y}_{k}),\hat{\phi}_{n}\rangle=2k+n,$
and let $\mathfrak{t}_{i}=\mathfrak{x}_{k}|Y_{n}(K)=\mathfrak{y}_{k}|Y_{n}(K)$.
@@
When $n\geq 2g(F)$, the chain complex $CF^{+}(Y_{n}(K),\mathfrak{t}_{i})$ is represented by the chain complex $A_{k,\mathfrak{t}}^{+}$, in the sense that there is an isomorphism
$\Psi_{n}^{+}:CF^{+}(Y_{n}(K),\mathfrak{t}_{i})\to A_{k,\mathfrak{t}}^{+}.$
Moreover, the squares in the diagrams of Figure 1 commute.
[…]
Figure 1: Commutative diagrams of the maps $v_{k}^{+}$ and $h_{k}^{+}$
The $CF^{\delta}$ version of the theorem holds with the same proof.
@@
${\mathbb{A}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,A_{\lfloor(i+pr)/q\rfloor,\mathfrak{t}}^{+}(K))\hskip 14.22636pt\text{and}\hskip 14.22636pt{\mathbb{B}}_{p/q,i,\mathfrak{t}}^{+}=\bigoplus_{r\in{\mathbb{Z}}}(r,B_{\mathfrak{t}}^{+}).$
@@
Let ${\mathbb{X}}_{p/q,i,\mathfrak{t}}^{+}$ denote the mapping cone of $D_{p/q,i,\mathfrak{t}}^{+}$; in section 4, we prove that the homology of this chain complex is isomorphic to $HF^{+}(Y_{p/q}(K),i,\mathfrak{t})$ [OS10].
@@ 3 Integer Surgeries
Fix $n>0$. The long exact sequence in Theorem 2.2 decomposes over $\operatorname{Spin}^{c}$ structures on $Y$. Setting $m=nk$ for large $k>0$, it also decomposes over $\operatorname{Spin}^{c}$ structures on $Y_{n}(K)$, i.e. given $i\in{\mathbb{Z}}/n{\mathbb{Z}}$, $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{+})$ where
$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$ (3.1)
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k+1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k+1)\equiv 2s\pmod{2n(k+1)}$,
$\text{and }s\equiv i\pmod{n}\}$,
$CF^{+}(Y_{n(k+1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{+}(Y_{n(k+1)}(K),s,\mathfrak{t}).$
A strengthening of Proposition 3.5 in [NI20] shows the following:
@@
Let $n>0,$ $n/2\leq k<n/2$, and $p=\max\{k+g(F),-2k+g(F)\}$. Suppose that $n\geq\max\{p,2g(F)\}$, $\mathfrak{s}\in\mathcal{X}_{k}\setminus\{\mathfrak{x}_{k},\mathfrak{y}_{k}\},$ then the map $F_{W_{n}^{\prime}(K),\mathfrak{s}}^{-}$ factorizes through the map
@@
This proposition is exactly the same as [NI20] Proposition 3.5 when $k\leq 0$, but when $k>0$, we improve the bound on the power of $U$ from n2kg(F)n-2k-g(F) to np=nkg(F)n-p=n-k-g(F). The proof of the improved bound is just checking that when $k>0$, the inequality
$\frac{l(l+1)n}{2}-lk-g\geq n-2|k|-g$
in Ni’s proof can be improved to
$\frac{l(l+1)n}{2}-lk-g\geq n-k-g.$
By the same proof, the proposition also holds for the $\delta$ version of Heegaard Floer homology for any $\delta\geq 0$.
@@
is quasi-isomorphic to $CF^{+}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
@@
We first prove the theorem for the $CF^{\delta}$ version of knot Floer homology for any $\delta\geq 0$, that is, the mapping cone ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$ is quasi-isomorphic to $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ as ${\mathbb{Z}}[U]$-modules.
Recall that $CF^{\delta}(Y_{n}(K),i,\mathfrak{t})$ is quasi-isomorphic to $MC(f_{2,i,\mathfrak{t}}^{\delta})$ as in Equation 3.1. Since $c_{1}(\mathfrak{t})$ is nontorsion, $H_{*}(A_{i}^{+})$ is a finitely generated abelian group. So $U^{q}|H_{*}(A_{i}^{+})=0$ when $q$ is greater than a constant $C_{1}$ independent of $n,k$. Hence, $U^{q}|H_{*}(A_{i}^{\delta})=0$ for $q>C_{1}$ for the same constant $C_{1}$ independent of $n,k,$ and $\delta$. By the $CF^{\delta}$ version of Theorem 2.7, this implies that $U^{q}|HF^{\delta}(Y_{n(k+1)}(K),i,\mathfrak{t})=0$ when $q>C_{1}$ and $n(k+1)>2g$. Choose $k$ such that $n(k+1)>2(g+\delta)$.
Consider a $\operatorname{Spin}^{c}$ structure $\mathfrak{s}_{s}\in\mathcal{X}_{i}$ for some $s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}$. Recall that we choose representatives
$\frac{n(k+1)}{2}\leq s<\frac{n(k+1)}{2}$
@@
$l\leq\frac{n(k+1)}{2}+g\leq n(k+1).$
Proposition 3.2 implies that $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$ when $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$.
Now for $s<0$, let $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then by conjugation invariance, there is a chain homotopy equivalence
$F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}\cong F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}.$